How to find the slope intercept form of the line that satisfies the given
contains the point (1,3) and is parallel to x + 5y = 9?

Answer by Gogonati(855)

Write the equation x+5y=9 in the slope intercept form y=-(1/5)x+9/5. Since parallel lines have equal slopes, the equation of our line will be:
y-3=-(1/5)(x-1), converting this equation in the standard form, we get:
5(y-3)=1-x <=> 5y-15=1-x <=> x+5y=16.